OpenStudy (anonymous):

Use the arc length formula and the given information to find s. Show your work for full credit. r = 20 ft θ = 19° s = ?

OpenStudy (anonymous):

Length of arc is given by s = rθ ( here θ should be written in radians) \[θ = 19^0 = 19 \times \frac{\pi}{180^0} = \frac{19\pi}{180^0}\] \[s= r \times \theta = 20 \times \frac{19\pi}{180^0} = \frac{19\pi}{9}= \frac{19 \times 3.14 }{9}\] \[= \frac{59.66 }{9} = 6.628 \approx 6.63 ft\] Thus length of arc = 6.63 ft @suzanna

OpenStudy (anonymous):

Oh wow thank you I was making sure I did it right and I did! Thank you so much!